Consider the following contingency table. 20 12 AC S0 a. Convert the contingency table into a joint probability table. (Round intermediate calculations to at least 4 declmal places and final answer to 4 decimal places.) Tota AC Total b. What is the probability that A occurs? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.) c. What is the probability that A and B occur? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.) Probability
d. Gven that B has occurred, final answer to 4 decimal places.) what is the probabity that A occurs? (Round intermediate calculations to at least 4 decimal places and Probability e. Given that A has occurred, what is the probability that Boccurs? (Round intermediate calculations to at least 4 decimal places and final answer to 4 decimal places.) Probability
Answer #1

dividing each cell with 100;

a)

B Bc total
A 0.2 0.12 0.32
Ac 0.5 0.18 0.68
total 0.7 0.3 1

b)

probability =0.32

c)

probability =0.2

d)

probability =20/70=0.2857

e)

probability =50/68=0.7353